OpenStudy (anonymous):

-15p^2w^8/9pw(7p^5) step by step please thanks

OpenStudy (angela210793):

-15p^2w^8/63p^6w=-15p^-4w^7/63=\[(-15w^7)/(63p^4)\]

OpenStudy (angela210793):

=(-5w^7)/(21p^4)

OpenStudy (anonymous):

how did you get to (-15w^7)/(63p^4) ?

OpenStudy (anonymous):

you there?

OpenStudy (angela210793):

yea i'm here well u see x^a/x^b=x^(a-b) tht's wht i used

OpenStudy (anonymous):

actually, how did you get-15w^7/63p^4 = -5w^7/21p^4 ? thanks (:

OpenStudy (anonymous):

i think you made a mistake its -5p^4/21w^7 right ?

OpenStudy (anonymous):

you there ?